OpenStudy (anonymous):

Solve this linear equation using integrating factors... Plz help me first arrange the equation in standard form.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ y }{ x }+2x+1\]

OpenStudy (anonymous):

Standard form: \[\frac{ dy }{ dx }-\frac{ y }{ x }=2x+1\]?

OpenStudy (anonymous):

I solved it